# A model-agnostic gravitational-wave background characterization algorithm

### v1

@@ VII Discussion and Conclusions
We perform three different representative injections and show their recoveries in Section V, to showcase the capabilities of our method in the context of GWB searches. We first inject a realistic BBH background, consistent with current BBH population models, employing a LIGO A+ noise curve. As a reminder, no individual signals were subtracted or removed. Increasing the observing time improves the degree to which we can recover injections, but nonetheless the A+ sensitivity will not be enough to resolve the high frequency turnover. Therefore, we turn to injections using CE sensitivity and find that, for all three example injections, we can consistently recover key features of the injected signals using this model.
@@
Next, we inject a smoothed BPL, which is a type of signal we expect to see from, e.g. FOPTs in the early universe. We are able to accurately recover the injected signal, and also use the flexible model to place posteriors on model parameters. In this example, we construct posteriors on the recovered turnover frequency of the injected smoothed BPL. This showcases the power of the flexible algorithm, which could be applied in many other cases. Additionally, we find that even if the model were given flexibility to choose the number of interpolating knots, it tends to minimize the necessary number of knots, reducing the prior volume dynamically, and eliminating the need to manually test a number of possible models of increasing complexity.
Finally, we employ our algorithm to directly infer the merger rate of BBHs as a function of redshift and use this to construct a model for $\Omega_{\textrm{GW}}(f)$. As shown in Section VI, we can recover the injected $\mathcal{R}(z)$ using our flexible model assuming CE sensitivity. However, in this case, the credible intervals on $\mathcal{R}(z)$ broaden considerably for $z\gtrsim 2.5$.
Although we have chosen a set of representative examples, there are numerous extensions and use cases for this method. While we choose to recover the merger rate as a function of the redshift, we could have chosen the mass distribution, for example. Future work could also include fitting in multiple dimensions instead of just a single dimension. For example, we could implement 2D splines to allow for joint constraints on mass distribution vs. redshift, or beyond. Population analyses of individual events have already been extended to several dimensions [119, 120]. This feature of our injection and recovery procedure is useful because, given a detected background signal, we could recover real astrophysical distributions.
@@ Acknowledgements.
We would like to thank Katerina Chatziioannou for insightful discussions about this work and Jandrie Rodriguez for helping with preliminary investigations for this work. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. This work was supported by the National Science Foundation Research Experience for Undergraduates program through NSF grant PHY-2150027, the LIGO Laboratory Summer Undergraduate Research Fellowship program, and the California Institute of Technology Student-Faculty Programs. The computations presented here were conducted partly in the Resnick High Performance Computing Center, a facility supported by Resnick Sustainability Institute at the California Institute of Technology.
@@ A.1 RJMCMC Algorithm
The acceptance probability of moving from model $\mathbf{m}$ with $n$ knots activated to model $\mathbf{m}^{\prime}$ with $n^{\prime}$ knots activated is given by:
@@ A.2 Priors
We consider an overall prior, $p(\mathbf{m})$. The number of knots that are in use is $n$, the amplitude of those knots is $\mathbf{a}_{n}$, and the location of the knots is $\mathbf{x}_{n}$. The prior is $p(\mathbf{a}_{n},\mathbf{x}_{n},n)$, which we expand using the product rule
@@
The prior probability of our amplitudes and horizontal positions are independent of one another. So we can expand our prior even further
@@
where the prior on $\mathbf{a}_{n}$ and $\mathbf{x}_{n}$ are independent of each other, and priors on individual knots are also independent of the prior on other knots.
We place a uniform prior on our amplitudes between points $a_{\mathrm{min}}$ and $a_{\mathrm{max}}$,
$p(\mathbf{a}_{n}|n)=\frac{1}{\left(a_{\mathrm{max}}-a_{\mathrm{min}}\right)^{n}}$ (19)
@@
Finally, we set a uniform prior on the number of activated control points
@@
[…]
@@ A.3.1 Birth Proposal
In this case, we propose to move from a model with $\mathbf{m}=(\mathbf{a}_{n},\mathbf{x}_{n},n)$ to a new model with an extra knot $\mathbf{m}^{\prime}=(\mathbf{a}_{n+1},\mathbf{x}_{n+1},n+1)$. We note that all knot locations $\mathbf{x}_{n}$, are the same in $\mathbf{x}_{n+1}$, except for the addition of one extra knot, $x^{\prime}$. We choose $x^{\prime}$ by choosing randomly from the $N-n$ points that are currently deactivated. We choose the new grid point $x^{\prime}_{j}$ uniformly within $x_{\alpha}\pm\Delta x$, where $x_{\alpha}$ is the corresponding center of the underlying grid point.
@@
$\frac{q(\mathbf{m}|\mathbf{m}^{\prime})}{q(\mathbf{m}^{\prime}|\mathbf{m})}=2\Delta x\frac{p_{d}}{p_{b}}\left[\frac{p_{u}}{a_{\mathrm{max}}-a_{\mathrm{min}}}-(1-p_{u})\frac{e^{-\frac{1}{2}\frac{(a^{\prime}-a_{m})^{2}}{2\sigma^{2}}}}{\sqrt{2\pi\sigma^{2}}}\right]^{-1},$ (25)
@@
$\mathbf{R}(\mathbf{m}^{\prime}|\mathbf{m})=\frac{p_{d}}{p_{b}}\left[p_{u}-(1-p_{u})\frac{(a_{\mathrm{max}}-a_{\mathrm{min}})e^{-\frac{1}{2}\frac{(a^{\prime}-a_{m})^{2}}{2\sigma^{2}}}}{\sqrt{2\pi\sigma^{2}}}\right]^{-1}.$ (26)
@@ A.3.2 Death proposal
The death proposal is conceptually similar to the birth proposal. We propose to move from the current model, $\mathbf{m}=(\mathbf{a}_{n},\mathbf{x}_{n},n)$ to a new model $\mathbf{m}^{\prime}=(\mathbf{a}_{n-1},\mathbf{x}_{n-1},n-1)$. We randomly deactivate one of the $n$ activated points, $x^{\prime}$. The probability of moving from $\mathbf{m}$ to $\mathbf{m}^{\prime}$ corresponds to the probability that we chose that specific $x^{\prime}$. Likewise, the reverse probability of moving from $\mathbf{m}^{\prime}$ back to $\mathbf{m}$ is the probability that we have chosen to reactivate $x^{\prime}$ from a choice of $N-n+1$ available deactivated points, multiplied by the probability of choosing $a^{\prime}$ and $x^{\prime}$. This must take into account both of the methods for choosing a new point that we discussed in the previous section. Putting these things together, we find
$\mathbf{R}(\mathbf{m}^{\prime}|\mathbf{m})=\frac{p_{b}}{p_{d}}\left[p_{u}-(1-p_{u})\frac{(a_{\mathrm{max}}-a_{\mathrm{min}})e^{-\frac{1}{2}\frac{(a^{\prime}-a_{m}^{\prime})^{2}}{2\sigma^{2}}}}{\sqrt{2\pi\sigma^{2}}}\right],$ (27)
@@ A.3.3 Other Proposals
We consider three other MCMC proposals, all of which have $\mathbf{R}=1$:
@@
    Redraw the $x$-value of a randomly chosen knot within its valid prior region.
@@ Appendix B Bayes factors and odds ratios
By allowing the dimension of the interpolation model to vary, we naturally probe a range of distinct models, each separate configuration conceivably representing a distinct model. The final set of posterior samples, when properly converged, represents a posterior across those models. We can use these samples to compare the preference for different models. In the simplest example, consider fitting $\Omega_{\textrm{GW}}(f)$–one could construct an odds ratio between a signal model ($n>0$) and a noise model ($n=0$) as follows:
$O_{\mathcal{N}}^{S}=\frac{p(n>0|d)}{p(n=0|d)}\approx\frac{\textrm{\# of samples with }n>0}{\textrm{\# of samples with }n=0}\,.$ (28)
If it is equally likely that the number of activated knots range from $n=0$ to $n=N_{\mathrm{max}}$, then $p(n=k)=1/(N_{\mathrm{max}}+1)$. This means that the prior odds for the signal model to the noise model is given by:
$\frac{p(n>0)}{p(n=0)}=\frac{N_{\mathrm{max}}/(N_{\mathrm{max}}+1)}{1/(N_{\mathrm{m}ax}+1)}=N_{\mathrm{m}ax}.$ (29)
This means, then, that in the simplest case, we put in prior odds that favor the presence of a signal compared to noise–which is not typically what we want when we are making a detection. If we want equal prior odds on signal and noise (i.e. a Bayes factor!). This means that the Bayes factor between signal and noise is given by:
$B_{\mathcal{N}}^{S}=\frac{1}{N_{\mathrm{max}}}O_{\mathcal{N}}^{S}\approx\frac{1}{N}\frac{\textrm{\# of samples with }n>0}{\textrm{\# of samples with }n=0}.$ (30)
However, the interpolation model with large $N_{\mathrm{max}}$ has a much larger prior volume than the typical PL model that is used. So we can explore the effect on model selection of the increased Occam penalty incurred by adding this larger prior volume.
Let’s consider a toy model with no signal present. We place a uniform prior on the number of knots $n$. If we consider the evidence for the signal model in a vacuum on its own, then we use $p(n)=1/N_{\mathrm{max}}$. The evidence for signal model is
@@
$=\sum_{n=1}^{N_{max}}$ $\left[\frac{1}{N_{\mathrm{max}}}\frac{1}{(a_{max}-a_{min})^{n}}\right.$ (32)
×$\times\left.\prod_{i}\int_{a_{min}}^{a_{max}}\mathcal{L}_{N}\Theta(a_{i}-a_{crit})da_{i}\right]\,.$ (33)
@@
We note that in the second line, we assume the amplitudes at different nodes are independent (likely a faulty assumption in the end, but this is a toy model), and that the likelihood is roughly constant, $\mathcal{L}_{N}$, up to some value $a_{crit}$ before sharply falling to zero. We model this with the Heaviside function $\Theta$. The prior on $a$ and $n$ have been pulled out. If we write $\xi=(a_{crit}-a_{min})/(a_{max}-a_{min})$, then we find:
$\mathcal{Z}_{S}$ $=\frac{\mathcal{L}_{N}}{N_{max}}\sum_{n=1}^{N_{max}}\xi^{n}\,,$ (34)
$=\frac{\mathcal{L}_{N}}{N_{\mathrm{m}ax}}\xi\frac{1-\xi^{N_{\mathrm{m}ax}}}{1-\xi}\textrm{ assuming $\xi<1$}.$ (35)
The odds ratio for signal-to-noise, then, is given by:
$\mathcal{O}_{\mathcal{N}}^{S}=\frac{\mathcal{Z}_{S}}{\mathcal{Z}_{\mathcal{N}}}\frac{p(n>0)}{p(n=0)}\approx\xi\frac{1-\xi^{N_{\mathrm{m}ax}}}{1-\xi}.$ (36)
@@
We show how this scales for different values of $\xi$ (colors) and $N_{\mathrm{max}}$ in Fig. 9. We can see that for $\xi\approx 1$, there is essentially no Occam penalty for adding a spline knot, and so the odds ratio even when no signal is present, can grow with $N_{\mathrm{max}}$. However, for $\xi\lesssim 0.5$, the odds ratio does not grow significantly with $N_{max}$.
@@
Figure 9: We show $\mathcal{O}^{S}_{\mathcal{N}}$ for our toy model when no signal is present, for different choices of $N_{\mathrm{max}}$ (x-axis) and $\xi$ (colors). We see that for $\xi\approx 1$ the odds grow with $N_{max}$. But for $\xi\lesssim 0.5$ they grow slower, and return odds that are indifferent to either model.
This ambiguity–the prior odds in one case (small $\xi$) offsetting the very large Occam penalty incurred by allowing for complex models, versus the alternative case where $\xi\approx 1$ and the odds grow with $N_{\mathrm{max}}$ is an unavoidable fact of our model. In the case where we are uncertain of a detection, we naturally expect low detectable complexity in our model, and so it would make sense to choose $N_{\mathrm{max}}$ small (e.g. $\lesssim 5$ or so). This still allows us to potentially be more sensitive to, e.g. a PL that breaks in our sensitive band, but keep the detection-power of a simple PL model.
@@
The simplest way to deal with this, which we recommend for future analyses where simulations are cheap, is to first choose $N_{max}$, then create many simulated frequency-domain datasets with $\Omega_{gw}(f)=0$ and a variance $\sigma^{2}(f)$, and calculate a distribution on $\mathcal{O}^{S}_{N}$. One can then calculate a false alarm rate for $\mathcal{O}^{S}_{N}$ on the real data, under this noise model. This eliminates the ambiguity introduced into the odds ratio (and therefore Bayes factors) due to the choices in maximum model complexity and prior bounds.
@@
Finally, one can also calculate Bayes factors by estimating the evidence for signal $\mathcal{Z}_{S}$, and noise $\mathcal{Z}_{\mathcal{N}}$, directly using nested sampling [121] or thermodynamic integration [122, 123]. This allows for arbitrarily high Bayes factors to be calculated, as opposed to being limited to Bayes factors as large as the number of MCMC samples used, and is what is used for methods like BayesWave [87, 88]. However, the problem of calibrating the Bayes factors persists.
Appendix C Linear v. spline and Akima interpolation

### v2

@@ VII Discussion and Conclusions
We perform three different representative injections and show their recoveries in Sec. V, to showcase the capabilities of our method in the context of GWB searches. We first inject a realistic BBH background, consistent with current BBH population models, employing a LIGO A+ noise curve. As a reminder, no individual signals were subtracted or removed. Increasing the observing time improves the degree to which we can recover injections, but nonetheless the A+ sensitivity will not be enough to resolve the high-frequency turnover. Therefore, we turn to injections using CE sensitivity and find that, for all three example injections, we can consistently recover key features of the injected signals using this model.
@@
Next, we inject a smoothed BPL, which is a type of signal we expect to see from, e.g., FOPTs in the early Universe. We are able to accurately recover the injected signal, and also use the flexible model to place posteriors on model parameters. In this example, we construct posteriors on the recovered turnover frequency of the injected smoothed BPL. This showcases the power of the flexible algorithm, which could be applied in many other cases. Additionally, we find that even if the model were given flexibility to choose the number of interpolating knots, it tends to minimize the necessary number of knots, reducing the prior volume dynamically, and eliminating the need to manually test a number of possible models of increasing complexity.
Finally, we employ our algorithm to directly infer the merger rate of BBHs as a function of redshift and use this to construct a model for $\Omega_{\textrm{GW}}(f)$. As shown in Sec. VI, we can recover the injected $\mathcal{R}(z)$ using our flexible model assuming CE sensitivity. However, in this case, the credible intervals on $\mathcal{R}(z)$ broaden considerably for $z\gtrsim 2.5$.
Although we have chosen a set of representative examples, there are numerous extensions and use cases for this method. While we choose to recover the merger rate as a function of the redshift, we could have chosen the mass distribution, for example. Future work could also include fitting in multiple dimensions instead of just a single dimension. For example, we could implement 2D splines to allow for joint constraints on mass distribution vs redshift, or beyond. Population analyses of individual events have already been extended to several dimensions [heinzelNonparametricAnalysisCorrelations2024, heinzelPixelPopHighResolution2024]. This feature of our injection and recovery procedure is useful because, given a detected background signal, we could recover real astrophysical distributions.
@@ Acknowledgements.
We thank Katerina Chatziioannou for insightful discussions about this work and Jandrie Rodriguez for helping with preliminary investigations for this work. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation (NSF). The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants No. PHY-0757058 and No. PHY-0823459. This work was supported by the National Science Foundation Research Experience for Undergraduates program through NSF Grant No. PHY-2150027, the LIGO Laboratory Summer Undergraduate Research Fellowship program, and the California Institute of Technology Student-Faculty Programs. The computations presented here were conducted partly in the Resnick High Performance Computing Center, a facility supported by Resnick Sustainability Institute at the California Institute of Technology.
@@ A.1 RJMCMC Algorithm
The acceptance probability of moving from model $\mathbf{m}$ with $n$ knots activated to model $\mathbf{m}^{\prime}$ with $n^{\prime}$ knots activated is given by
@@ A.2 Priors
We consider an overall prior $p(\mathbf{m})$. The number of knots that are in use is $n$, the amplitude of those knots is $\mathbf{a}_{n}$, and the location of the knots is $\mathbf{x}_{n}$. The prior is $p(\mathbf{a}_{n},\mathbf{x}_{n},n)$, which we expand using the product rule
@@
The prior probability of our amplitudes and horizontal positions are independent of one another. So we can expand our prior even further:
@@
where the prior on $\mathbf{a}_{n}$ and $\mathbf{x}_{n}$ are independent of each other and priors on individual knots are also independent of the prior on other knots.
We place a uniform prior on our amplitudes between points $a_{\textrm{min}}$ and $a_{\textrm{max}}$:
$p(\mathbf{a}_{n}|n)=\frac{1}{\left(a_{\textrm{max}}-a_{\textrm{min}}\right)^{n}}$ (19)
@@
Finally, we set a uniform prior on the number of activated control points:
@@
$\sum_{n=1}^{N}\left\{p(n)\left[\prod_{j}\int dx_{j}\,p(x_{j}|n)\right]\left[\prod_{j}\int_{a_{\textrm{min}}}^{a_{\textrm{max}}}\,da_{j}\;p(a_{j}|n)\right]\right\}=1.$ (22)
@@ A.3.1 Birth proposal
In this case, we propose to move from a model with $\mathbf{m}=(\mathbf{a}_{n},\mathbf{x}_{n},n)$ to a new model with an extra knot $\mathbf{m}^{\prime}=(\mathbf{a}_{n+1},\mathbf{x}_{n+1},n+1)$. We note that all knot locations $\mathbf{x}_{n}$, are the same in $\mathbf{x}_{n+1}$, except for the addition of one extra knot, $x^{\prime}$. We choose $x^{\prime}$ by choosing randomly from the NnN-n points that are currently deactivated. We choose the new grid point $x^{\prime}_{j}$ uniformly within $x_{\alpha}\pm\Delta x$, where $x_{\alpha}$ is the corresponding center of the underlying grid point.
@@
$\frac{q(\mathbf{m}|\mathbf{m}^{\prime})}{q(\mathbf{m}^{\prime}|\mathbf{m})}=2\Delta x\frac{p_{d}}{p_{b}}\left[\frac{p_{u}}{a_{\textrm{max}}-a_{\textrm{min}}}-(1-p_{u})\frac{e^{-\frac{1}{2}\frac{(a^{\prime}-a_{m})^{2}}{2\sigma^{2}}}}{\sqrt{2\pi\sigma^{2}}}\right]^{-1},$ (25)
@@
$\mathbf{R}(\mathbf{m}^{\prime}|\mathbf{m})=\frac{p_{d}}{p_{b}}\left[p_{u}-(1-p_{u})\frac{(a_{\textrm{max}}-a_{\textrm{min}})e^{-\frac{1}{2}\frac{(a^{\prime}-a_{m})^{2}}{2\sigma^{2}}}}{\sqrt{2\pi\sigma^{2}}}\right]^{-1}.$ (26)
@@ A.3.2 Death proposal
The death proposal is conceptually similar to the birth proposal. We propose to move from the current model, $\mathbf{m}=(\mathbf{a}_{n},\mathbf{x}_{n},n)$ to a new model $\mathbf{m}^{\prime}=(\mathbf{a}_{n-1},\mathbf{x}_{n-1},n-1)$. We randomly deactivate one of the $n$ activated points, $x^{\prime}$. The probability of moving from $\mathbf{m}$ to $\mathbf{m}^{\prime}$ corresponds to the probability that we chose that specific $x^{\prime}$. Likewise, the reverse probability of moving from $\mathbf{m}^{\prime}$ back to $\mathbf{m}$ is the probability that we have chosen to reactivate $x^{\prime}$ from a choice of Nn+1N-n+1 available deactivated points, multiplied by the probability of choosing $a^{\prime}$ and $x^{\prime}$. This must take into account both of the methods for choosing a new point that we discussed in the previous section. Putting these things together, we find
$\mathbf{R}(\mathbf{m}^{\prime}|\mathbf{m})=\frac{p_{b}}{p_{d}}\left[p_{u}-(1-p_{u})\frac{(a_{\textrm{max}}-a_{\textrm{min}})e^{-\frac{1}{2}\frac{(a^{\prime}-a_{m}^{\prime})^{2}}{2\sigma^{2}}}}{\sqrt{2\pi\sigma^{2}}}\right],$ (27)
@@ A.3.3 Other proposals
We consider three other MCMC proposals, all of which have $\mathbf{R}=1$.
@@
    Redraw the $x$ value of a randomly chosen knot within its valid prior region.
@@ Appendix B BAYES FACTORS AND ODDS RATIOS
By allowing the dimension of the interpolation model to vary, we naturally probe a range of distinct models, where each configuration conceivably represents a different model. The final set of posterior samples, when properly converged, represents a posterior across those models. We can use these samples to compare the preference for different models. However, the transdimensional nature of the comparison leads to irreducible ambiguities when defining a metric for model selection–e.g., Bayes’ factors. We elaborate on this here, and provide suggestions for Bayes’ factor calibration.
Consider fitting $\Omega_{\textrm{GW}}(f)$; one could construct an odds ratio between a signal model ($n>0$) and a noise model ($n=0$) as follows:
@@
If it is equally likely that the number of activated knots range from $n=0$ to $n=N_{\textrm{max}}$, then $p(n=k)=1/(N_{\textrm{max}}+1)$. This means that the prior odds for the signal model to the noise model is given by:
$\frac{p(n>0)}{p(n=0)}=\frac{N_{\textrm{max}}/(N_{\textrm{max}}+1)}{1/(N_{\textrm{m}ax}+1)}=N_{\textrm{max}}.$ (29)
In this case we have designed prior odds that favor the presence of a signal compared to noise. This is not a typical choice in the context of signal detection, where it is often more preferable to have false negative results compared to false positive results. A statistic with equal prior odds on signal and noise (a Bayes factor) is more common and is given by
[…]
. Conversely, the interpolation model with large $N_{\textrm{max}}$ has a much larger prior volume than the typical PL model that is used. This is not a problem if the extra parameters fill their prior distributions. However in stochastic analyses like those presented here, very broad uninformative priors are often used, such that the data produce upper limits well below the prior that was set. This can result in unintended outcomes.
To explore the push and pull between the prior odds and the Occam penalty for our specific case more concretely, we consider a representative toy example that captures the spirit of the problem. We place a uniform prior on the number of knots $n$. If we consider the evidence for the signal model only, then we use $p(n)=1/N_{\textrm{max}}$. The evidence for the signal model is then
$\mathcal{Z}_{S}=\sum_{n=1}^{N_{max}}$ $\int p(d|a,n)p(a,n)da\,$, (31)
$=\sum_{n=1}^{N_{max}}$ $\left[\frac{1}{N_{\textrm{max}}}\frac{1}{(a_{max}-a_{min})^{n}}\right$. (32)
$\times\left.\prod_{i}\int_{a_{min}}^{a_{max}}\mathcal{L}_{N}\Theta(a_{i}-a_{crit})da_{i}\right]\,.$ (33)
@@
In Eq.  (31), we assume the amplitudes at different nodes are independent (likely a faulty assumption in the end, but this is a toy example) and that the likelihood is roughly constant and equal to $\mathcal{L}_{N}$, up to some value $a_{crit}$ before sharply falling to zero. This is the typical situation for GWB analyses upper limits are set on the GWB can be set as a function of frequency. We model this with the Heaviside function $\Theta$. The priors on $a$ and $n$ have been pulled out. Defining $\xi=(a_{crit}-a_{min})/(a_{max}-a_{min})$, the ratio between the size of the posterior and the size of the prior, we find:
$\mathcal{Z}_{S}$ $=\frac{\mathcal{L}_{N}}{N_{\textrm{max}}}\sum_{n=1}^{N_{max}}\xi^{n}\,$, (34)
$=\frac{\mathcal{L}_{N}}{N_{\textrm{max}}}\xi\frac{1-\xi^{N_{\textrm{max}}}}{1-\xi}\textrm{ assuming $\xi<1$}.$ (35)
The odds ratio for signal to noise, then, is given by
$\mathcal{O}_{\mathcal{N}}^{S}=\frac{\mathcal{Z}_{S}}{\mathcal{Z}_{\mathcal{N}}}\frac{p(n>0)}{p(n=0)}\approx\xi\frac{1-\xi^{N_{\textrm{m}ax}}}{1-\xi}.$ (36)
@@
We show how this scales for different values of $\xi$ (colors) and $N_{\textrm{max}}$ in Fig. 10. We can see that for $\xi\approx 1$, there is essentially no Occam penalty for adding a spline knot, and so the odds ratio even when no signal is present, can grow with $N_{\textrm{max}}$. However, for $\xi\lesssim 0.5$, the odds ratio does not grow significantly with $N_{\rm max}$.
@@
Figure 10: We show $\mathcal{O}^{S}_{\mathcal{N}}$ for our toy model when no signal is present, for different choices of $N_{\textrm{max}}$ (x axis) and $\xi$ (colors). We see that for $\xi\approx 1$ the odds grow with $N_{\rm max}$. But for $\xi\lesssim 0.5$ they grow slower and return odds that are indifferent to either model.
This leads to an unavoidable ambiguity of our model: when $\xi$ is small, the prior odds offset the very large Occam penalty incurred by allowing for complex models, while when $\xi\approx 1$ the odds grow with $N_{\textrm{max}}$.
This implies that the transdimensional approach should be handled with care in the context of signal detection. When claiming a detection with a threshold significance, we expect low detectable complexity in our model, and so we advise choosing a small $N_{\textrm{max}}$, e.g. $\lesssim 5$. This still allows us to potentially be more sensitive to e.g. a PL that breaks in our sensitive band, but keep the detection power of a simple PL model. In this framework, frequentist statistics such as the SNR of Fig. 2 should provide a useful guideline.
@@
If the noise is well understood, one can employ a simulation-based approach to calibrate the odds ratio. This would involve first selecting an $N_{\rm max}$ and then creating many simulated frequency-domain datasets with $\Omega_{\textrm{GW}}(f)=0$ and a variance $\sigma^{2}(f)$ and calculating a distribution on $\mathcal{O}^{S}_{N}$. One can then calculate a false alarm rate for $\mathcal{O}^{S}_{N}$ on the real data, under this noise model. This would eliminate the ambiguity introduced into the odds ratio (and therefore Bayes factors) due to the choices in maximum model complexity and prior bounds.
@@
Finally, one can also calculate Bayes factors by estimating the evidence for signal $\mathcal{Z}_{S}$, and noise $\mathcal{Z}_{\mathcal{N}}$, directly using nested sampling [skilling_nested_2006] or thermodynamic integration [gelmanSimulatingNormalizingConstants1998, lartillotComputingBayesFactors2006a]. This allows for arbitrarily high Bayes factors and odds ratios to be calculated, while the RJMCMC technique is limited to odds ratios as large as the number of MCMC samples used. This is the method employed by BayesWave [Cornish:2014kda, Cornish:2020dwh].
Appendix C LINEAR VS SPLINE AND AKIMA INTERPOLATION